# IMPACT[Uncaptioned image]: Intelligent Motion Planning with Acceptable Contact Trajectories via Vision-Language Models

### v1

@@ Abstract
Motion planning involves determining a sequence of robot configurations to reach a desired pose, subject to movement and safety constraints. Traditional motion planning finds collision-free paths, but this is overly restrictive in clutter, where it may not be possible for a robot to accomplish a task without contact. In addition, contacts range from relatively benign (e.g., brushing a soft pillow) to more dangerous (e.g., toppling a glass vase). Due to this diversity, it is difficult to characterize which contacts may be acceptable or unacceptable. In this paper, we propose IMPACT, a novel motion planning framework that uses Vision-Language Models (VLMs) to infer environment semantics, identifying which parts of the environment can best tolerate contact based on object properties and locations. Our approach uses the VLM’s outputs to produce a dense 3D “cost map” that encodes contact tolerances and seamlessly integrates with standard motion planners. We perform experiments using 20 simulation and 10 real-world scenes and assess using task success rate, object displacements, and feedback from human evaluators. Our results over 3620 simulation and 200 real-world trials suggest that IMPACT enables efficient contact-rich motion planning in cluttered settings while outperforming alternative methods and ablations. Supplementary material is available at https://impact-planning.github.io/.
@@ I Introduction
Classical motion planning for robot manipulation [1, 2] frames the problem as finding a path for the robot’s end-effector to reach a target while avoiding collisions with obstacles. This formulation is generally desirable, but can be highly restrictive, especially in densely cluttered environments. In such cases, some incidental contact may be necessary to achieve a task, or to accomplish it more efficiently than by strictly avoiding all collisions.
@@
Consider the motivating example in Fig. 1, which shows a robot manipulator making contact with a toy bear to efficiently reach the target salt shaker. Due to the clutter, a collision-free path to the salt shaker either does not exist or would require a longer parabolic motion to go above the obstacles (which, again, may not be feasible in cluttered cabinets and boxes). We thus desire robots that can achieve a task while moving through a cluttered environment, making appropriate contact as needed. In this work, the term “contacts” does not refer to when a robot uses its grippers to touch a target object (e.g., for grasping). Instead, we consider other types of contact: when any part of the robot touches any non-target (or “distractor”) object in the environment. We study this in motion planning for tasks that involve reaching to a target in dense clutter.
To address this challenge, we propose Intelligent Motion Planning with Acceptable Contact Trajectories (IMPACT), a novel framework that leverages modern Vision-Language Models (VLMs) such as GPT-4o [3] to infer object contact tolerances. After analyzing object properties, IMPACT generates a 3D cost map to represent the tolerance rate of different regions in the scene. This cost map integrates seamlessly with standard off-the-shelf motion planning algorithms to enable trajectories which engage in efficient and semantically-acceptable contact. We pair IMPACT with the RRT* motion planning algorithm [4] as our primary method, and we refer to this combination as IMPACT+RRT*. We perform experiments in simulation and real-world settings, testing a variety of contact-rich, densely cluttered reaching tasks. We evaluate IMPACT+RRT* using multiple quantitative metrics, including human preference rankings from a user study. Our results indicate that users prefer IMPACT+RRT* over alternatives, suggesting that it is a promising approach for facilitating semantically-acceptable contact-rich manipulation.
@@
Figure 1: An example of a reaching task and object costs. The first row shows the difference between collision-free paths and paths with acceptable contact. Left: Collision-free paths prevent a “straight” path to the salt shaker because of the toy bear and the glass bottle obstacles (each marked with a red “X”). Right: With semantically acceptable contact, the robot can successfully reach the salt shaker by pushing the toy bear and avoiding the fragile glass bottle. The second row shows the cost of each object generated by GPT-4o. Left: the original scene. Right: GPT-4o assigns different costs to objects, with the target assigned -1 (toy bear: 4, salt shaker: -1, glass bottle: 9).
@@
  • IMPACT, a framework that formalizes “acceptable contact” and generates a 3D cost map to densely represent object tolerance information.
  • A motion planning pipeline that integrates contact cost maps with standard motion planning to find semantically-acceptable contact-rich trajectories.
  • Simulation and real-world experiments of IMPACT+RRT* for reaching-based tasks in dense clutter, and evaluation of contact-rich trajectories using quantitative task-based and human feedback metrics.
@@ II-A Robot Motion Planning
Classical motion planning algorithms can be broadly characterized as optimization-based or sampling-based. Optimization-based methods, such as TrajOpt [5], frame the problem as minimizing a cost function subject to kinematic, dynamics, and obstacle avoidance constraints. Sampling-based methods, such as PRM [6], RRT [7], and RRT* [4], incrementally build a graph or tree of feasible paths. While there are numerous variants of these methods, a common underlying theme is the constraint of avoiding any collisions. This is generally desirable, but means motion planning under this standard formulation has limitations. In densely cluttered scenarios where contact is inevitable, these motion planners might not find a solution, even though one might exist if the robot engages in light contact. Our method seamlessly integrates with prior motion planning algorithms for flexible contact-rich manipulation.
Some motion planning works modify cost functions so that certain obstacles, such as leaves, are permeable [8, 9] and thus allow for some contact. However, these methods have been in domain-specific foliage settings and not tested in common cluttered scenarios with rigid objects. In closely-related work, [10] proposes to use semantic language commands to enable semantically-acceptable contact. Unlike [10], we do not require explicit language instructions about which contacts are acceptable or not, since we utilize VLMs to automate this process.
@@ II-C Vision-Language Models (VLMs) for Robotics
VLMs such as GPT-4o [3] and Gemini [17] are trained on broad Internet-scale data and have remarkable semantic and spatial knowledge. Thus, the research community has explored numerous applications of VLMs in robotics [18, 19, 20]. One way to use VLMs in robotics is to generate high-level task plans, in the form of natural language [21, 22, 23] or executable code [24, 25]. Another way to use VLMs is for reward [26, 27, 28] or task design [29, 30, 31]. VLMs can also aid low-level affordance reasoning [32, 33] which, in an extension of this direction, can involve generating full trajectories [34]. In contrast to these works, we use VLMs for a complementary objective: to infer object contact tolerances to guide contact-rich motion planning.
In closely-related work, VoxPoser [35] exploits VLMs for open-world reasoning and visual grounding to compose a 3D value map to guide robotic interactions. As in [35], we use VLMs to construct a 3D value map (equivalently, a cost map). However, unlike VoxPoser, which requires the user to explicitly specify which objects to avoid with language, e.g., “get the item, but watch out for that vase!” we do not require explicit language commands. Instead, we leverage the improving spatial and semantic knowledge of recent VLMs [36] to determine object contact tolerances, and show manipulation in more densely cluttered environments. Other recent work relies on VLMs for semantically-safe manipulation [37] but assumes that any collisions are undesirable.
@@ III Problem Statement and Assumptions
We assume a single robot arm with a standard gripper operates in a densely cluttered environment that contains $n$ objects, denoted as ${\mathcal{O}=\{o_{1},o_{2},\ldots,o_{n}\}}$. The robot must reach a given target object ${o_{\rm targ}\in\mathcal{O}}$ while minimizing unwanted contact with other objects (i.e., obstacles) in ${\mathcal{O}\setminus o_{\rm targ}}$. In environments with significant clutter, $o_{\rm targ}$ may be behind or close to multiple objects and thus reaching it may be infeasible with a collision-free trajectory. To reduce occlusions, at least two cameras provide respective RGBD images at the start of the task. Given these image observations, the objective is to compute a trajectory $\tau$ for the robot, defined as a sequence of gripper poses, such that its gripper ultimately touches $o_{\rm targ}$. Thus, among the dynamically feasible trajectories, our objective is to select one that reaches the target while engaging in “semantically-acceptable” contact with obstacles when needed.
@@ IV Method: IMPACT
Figure 2: Overview of IMPACT. There is a toy bear, a coffee cup and a tomato can on the table. The objective is to reach the tomato can. We use SAM2 [38] to segment the image and label the objects using “1,” “2,” and “3” to assist GPT’s visual reasoning. GPT also receives a language template prompt $\ell$ with object information from SAM2. GPT produces costs for the three objects, which are projected into a voxel grid $C$ indicating the cost for the robot end-effector to enter each voxel. The costs are high for the coffee cup (GPT-assigned cost: 8) and the tabletop (a fixed cost of 10). We use a cost of -1 for the target object. Finally, an off-the-shelf motion planner (RRT*) uses this to guide the robot, which avoids the coffee cup but makes contact with the toy bear to successfully reach the tomato can.
@@
IV-A Obtaining Object Costs using GPT
@@
We use GPT-4o (hereafter, GPT) as the VLM to generate the cost of all objects in the scene due to its strong spatial reasoning capabilities [36], but our approach is compatible with other VLMs. The input to GPT includes both an annotated image and a text prompt. To get the image input, we mount an RGBD camera between the scene and the robot to capture a front-view image $I$ in simulation, while in the real world setup, we mount a camera at the right front of the scene to get $I$. Cluttered scenes may contain too many objects for GPT to accurately recognize all of them. Consequently, we use SAM2 [38] to segment the objects. Then, we use Set-of-Mark prompting [39] to annotate the image with numerical indices over the objects. The segmented and annotated image $I^{\prime}$ is ultimately provided as part of the input to GPT.
@@
We also design a text prompt template $\ell$ which includes the list of objects in the scene (but is otherwise task-agnostic), and some general principles related to the concept of contact tolerance. Objects are labeled with the same numbers as in the input image, marked by SAM2. We include Chain-of-Thought prompting [40] in $\ell$ to improve GPT’s reasoning. The full text prompt can be found on our website. The output of GPT is a dictionary with the cost of all queried objects. Fig. 1 (second row) shows an example of generated object costs, suggesting that GPT can accurately reason about object properties. For example, it understands that a glass bottle is fragile, and thus assigns it a high cost of 9. In contrast, other objects have a lower cost (i.e., better contact tolerance) such as the toy bear with a cost of 4.
IV-B Motion Planning with Contacts
We use motion planning algorithms to synthesize robot trajectories in the cluttered environment. Inspired by the 3D value map in VoxPoser [35], we propose to construct a 3D cost map $C$ to guide the motion planner. The cost map is represented as a voxel grid of dimension $(L\times W\times H)$, where each voxel $C[x,y,z]$ denotes the cost at position $(x,y,z)$. We initialize $C$ to be all zero. Then, for each object in the scene, we identify the occupied voxels in 3D space and assign them the corresponding object cost from GPT (see Sec. IV-A). During planning, we further assign the cost of the target to -1 instead of using the cost generated by GPT. This encourages the planning algorithm to find a path towards the target. We visualize cost maps in Fig. 2 and Fig. 3.
Once we construct the 3D cost map, we use a standard motion planning algorithm to compute a trajectory that minimizes cost. Each trajectory consists of a sequence of robot end-effector positions. The total cost of a trajectory $\tau$ is the sum of costs of obstacles the robot arm collides with:
$Cost(\tau)=\sum_{i=1}^{|\tau|}C[\tau_{i}].$ (1)
In Equation 1, $|\tau|$ denotes the length of the trajectory $\tau$ and $\tau_{i}$ is the position of the $i$-th waypoint. By combining the cost map and the motion planning algorithm, we can compute a trajectory that minimizes contact with high-cost obstacles while permitting acceptable contact when necessary. Notably, we only use Equation 1 during motion planning. For evaluation, we also compute the cost based on the obstacles the robot arm makes contact with, but we do not count the same obstacle multiple times (see Sec. V-D).
As our main method, we use IMPACT with the RRT* [4] motion planner, and denote this as IMPACT+RRT*.
Figure 3: Our simulation setup in PyBullet simulation [41]. In this example scene, there is a box which contains a wine glass, a sugar box, and a mug. The objective is to reach the mug. The yellow and green regions (e.g., for the wine glass and the larger box) indicate higher costs. For visual clarity, we show the original scene $I$, before it is annotated for GPT. See Sec. V for more details.

### v2

@@ Abstract
Motion planning involves determining a sequence of robot configurations to reach a desired pose, subject to movement and safety constraints. Traditional motion planning finds collision-free paths, but this is overly restrictive in clutter, where it may not be possible for a robot to accomplish a task without contact. In addition, contacts range from relatively benign (e.g., brushing a soft pillow) to more dangerous (e.g., toppling a glass vase), making it difficult to characterize which may be acceptable. In this paper, we propose IMPACT, a novel motion planning framework that uses Vision-Language Models (VLMs) to infer environment semantics, identifying which parts of the environment can best tolerate contact based on object properties and locations. Our approach generates an anisotropic cost map that encodes directional push safety. We pair this map with a contact-aware A* planner to find stable contact-rich paths. We perform experiments using 20 simulation and 10 real-world scenes and assess using task success rate, object displacements, and feedback from human evaluators. Our results over 3200 simulation and 200 real-world trials suggest that IMPACT enables efficient contact-rich motion planning in cluttered settings while outperforming alternative methods and ablations. Our project website is available at https://impact-planning.github.io/.
@@ I Introduction
Classical motion planning for robot manipulation [25, 37] frames the problem as finding a path for the robot’s end-effector to reach a target while avoiding collisions with obstacles. This formulation is generally desirable, but can be highly restrictive, especially in densely cluttered environments [7]. In such cases, some incidental contact may be necessary to achieve a task, or to accomplish it more efficiently than by strictly avoiding all collisions.
@@
Consider the example in Fig. 1, which shows a robot manipulator making contact with a toy bear to efficiently reach the target spice jar. Due to the clutter, a collision-free path either does not exist or would require a longer parabolic motion to go above the obstacles (which, again, may not be feasible in cluttered cabinets and boxes). We thus desire robots that can achieve a task while moving through a cluttered environment, making appropriate contact as needed. In this work, the term “contacts” does not refer to when a robot uses its grippers to touch a target object (e.g., for grasping). Instead, we consider other types of contact: when any part of the robot touches any non-target (or “distractor”) object in the environment. We study this in motion planning for tasks that involve reaching to a target in dense clutter.
@@
Figure 1: An example of a reaching task and object costs. The first row shows the difference between collision-free paths and paths with acceptable contact. Left: Collision-free paths prevent a “straight” path to the spice jar because of the toy bear and wine glass obstacles (each marked with a red “X”). Right: With semantically acceptable contact, the robot can successfully reach the spice jar by pushing the toy bear and avoiding the fragile wine glass. The second row shows the cost of each object generated by GPT-4o. Left: the original scene. Right: GPT-4o assigns different costs to objects, with the target assigned 1-1 (toy bear: 33, spice jar: 1-1, wine glass: 88).
To address this challenge, we propose Intelligent Motion Planning with Acceptable Contact Trajectories (IMPACT), a novel framework that leverages modern Vision-Language Models (VLMs) such as GPT-4o [29] to infer initial semantic costs for objects in the scene. However, knowing that an object is generally safe to touch is different from knowing how and where to push it without causing a cascade of collisions. Our method repeatedly samples push outcomes to estimate the probability of causing a collision with obstacles or the target. This risk is condensed into safety scores that are used to construct an anisotropic cost map, where interaction costs vary with direction since pushing an object from one side versus another can lead to different outcomes. We integrate this directional cost map with a contact-aware A* planner [11] that computes paths to reach targets. We perform experiments on contact-rich, densely cluttered reaching tasks in simulation and real-world settings. We evaluate IMPACT using multiple quantitative metrics, including human preference rankings from a user study. Results indicate that users prefer IMPACT over alternatives, suggesting that it is a promising approach for facilitating semantically-acceptable contact-rich manipulation.
@@
  • IMPACT, a framework that formalizes “acceptable contact” by transforming initial VLM-inferred semantic costs into a dense, anisotropic cost map to represent the directional safety of physical interactions.
  • A contact-aware A* planner capable of interpreting this anisotropic map to execute paths with intelligent and minimal-impact contact.
  • Extensive simulation and real-world experiments, including a human subjects study, that validate our approach using both objective and subjective metrics to evaluate the success of contact-rich trajectories.
@@ II-A Robot Motion Planning
Classical motion planning algorithms can be broadly characterized as optimization-based or sampling-based [37]. Optimization-based methods, such as TrajOpt [34], frame the problem as minimizing a cost function subject to kinematic, dynamics, and obstacle avoidance constraints. Sampling-based methods, such as PRM [15], RRT [20], and RRT* [14], incrementally build a graph or tree of feasible paths. While there are numerous variants of these methods, a common underlying theme is the constraint of avoiding any collisions. This is generally desirable, but has limitations. In densely cluttered scenarios where contact is inevitable, these motion planners might not find a solution, even though one might exist if the robot engages in light contact. Our method seamlessly integrates with prior motion planning algorithms for flexible contact-rich manipulation.
Some motion planning works modify cost functions so that certain obstacles, such as leaves in foliage, are permeable [17, 28] and thus allow for some contact. These methods have been applied to domain-specific foliage settings and not tested in common cluttered scenarios with rigid objects. In closely-related work, Xie et al. [41] propose to use semantic language commands to enable semantically-acceptable contact. Unlike [41], we do not require explicit language instructions about which contacts are acceptable or not, since we utilize VLMs to automate this process.
@@ II-C Vision-Language Models (VLMs) for Robotics
VLMs such as GPT-4o [29] and Gemini [10] are trained on broad Internet-scale data and have remarkable semantic and spatial knowledge. Thus, the research community has explored numerous applications of VLMs in robotics [9, 16]. One way to use VLMs in robotics is to generate high-level task plans, in the form of natural language [1, 8] or executable code [23, 35]. Another way to use VLMs is for reward [26, 36] or task design [38, 39]. VLMs can also aid low-level affordance reasoning [18, 24] which, in an extension of this direction, can involve generating full trajectories [19]. In contrast to these works, we use VLMs for a complementary objective: to infer object contact tolerances to guide contact-rich motion planning.
In closely-related work, VoxPoser [12] exploits VLMs for open-world reasoning and visual grounding to compose a 3D value map to guide robotic interactions. As in [12], we use VLMs to assign semantic costs to objects. However, unlike VoxPoser, which requires the user to explicitly specify which objects to avoid with language, e.g., “get the item, but watch out for that vase!” we do not require explicit language commands. Instead, we leverage the improving spatial and semantic knowledge of recent VLMs [5] to determine object contact tolerances, and show manipulation in more densely cluttered environments. Other recent work relies on VLMs for semantically-safe manipulation [2] but assumes that any collisions are undesirable.
@@ III Problem Statement and Assumptions
We assume a single robot arm with a standard gripper operates in a densely cluttered environment that contains $n$ objects, denoted as $\mathcal{O}=\{o_{1},o_{2},\ldots,o_{n}\}$}. The robot must reach a given target object $o_{\rm targ}\in\mathcal{O}$} while minimizing unwanted contact with other objects (i.e., obstacles) in $\mathcal{O}\setminus o_{\rm targ}$}. In environments with significant clutter, $o_{\rm targ}$ may be behind or close to multiple objects and thus reaching it may be infeasible with a collision-free trajectory. To reduce occlusions, at least two cameras provide respective RGBD images at the start of the task. Given these image observations, the objective is to compute a trajectory $\tau$ for the robot, defined as a sequence of end-effector poses, such that its end-effector ultimately touches $o_{\rm targ}$. Thus, among the dynamically feasible trajectories, our objective is to select one that reaches the target while engaging in “semantically-acceptable” contact with obstacles when needed.
@@ IV Method: IMPACT
Figure 2: Overview of IMPACT. There is a toy bear, a coffee cup and a glue bottle (target) on the table. The VLM receives an annotated image $I^{\prime}$ and a language template prompt $\ell$ with object information from SAM2 [31], and outputs costs for the three objects. We use a cost of 1-1 for the target object. We construct a 3D voxel grid $V$ using these costs and then flatten it to produce an anisotropic, contact-aware cost map $M^{\prime}$. The contact-aware A* planner searches over three motion primitives in this map: Move, Push and Rotate to generate a trajectory. The planner’s state space includes the robot’s end-effector pose and the displaced positions of low-cost objects. These guide the robot to avoid the coffee cup but make contact with the toy bear at the appropriate direction to reach the glue bottle.
@@
IV-A Obtaining Object Costs using a VLM
@@
We use GPT-4o (hereafter, GPT) as the VLM to generate the cost of all objects in the scene due to its strong spatial reasoning capabilities [5], but our approach is compatible with other VLMs. The input to the VLM includes both an annotated image and a text prompt. We mount an RGBD camera to capture a scene image $I$, and use SAM2 [31] to segment and annotate the objects. The segmented and annotated image $I^{\prime}$ is ultimately provided as part of the input to GPT. We also design a text prompt template $\ell$ which includes the list of objects in the scene (but is otherwise task-agnostic), and some general principles related to the concept of contact tolerance. Objects are labeled with the same numbers as in the input image, marked by SAM2. The full text prompt can be found on our website. The output of GPT is a dictionary with the cost of all queried objects. Fig. 1 (second row) shows an example of object costs generated from GPT. It assigns a high cost of $8$ to the wine glass and a lower cost of $3$ to the toy bear, which indicates different tolerances to contact across objects.
IV-B Directional Contact-Aware Motion Planning
We use motion planning algorithms to synthesize robot trajectories. We construct a 3D voxel grid $V$ where each voxel $V[x,y,z]$ denotes the cost at position $(x,y,z)$. During planning, we assign all voxels corresponding to the target object a cost of 1-1 to encourage the planner to find a path to it. To evaluate directional contact information, we project $V$ into a 2D grid $M$ using a top-down view, where $M[x,y]=\max_{z}V[x,y,z]$}. From this 2D perspective, all obstacles are divided into low-cost (5 or lower) and high-cost (6 or higher) sets based on their VLM-assigned costs. To simplify planning, we define three motion primitives in 2D space: (i) Move, which translates the robot to one of eight neighboring positions; (ii) Rotate, which discretizes the end-effector orientation changes within $\pm 45^{\circ}$; and (iii) Push, which moves the end-effector within a bounded radius while making contact with an object. We assume the first two change the end-effector’s pose without affecting the environment. In contrast, the Push primitive denotes end-effector translations that involve some robot-object contact.
IV-B1 Anisotropic Cost Map Generation
The map $M$ encodes object costs independent of motion direction. In clutter, however, contact safety depends strongly on the direction where the robot approaches an object. To address this, we construct a new anisotropic cost map $M^{\prime}$ that augments each point with directional safety information. Defining $M^{\prime}$ over both 2D positions and 2D directions would be high-dimensional and computationally expensive, so we keep $M^{\prime}$ as a 2D function. For each point $(x,y)$ on the boundary of a low-cost object, we consider the reverse surface normal vector, which points “inwards” towards the object. This gives the most natural and direct push for the object, but to consider different possible pushes, we sample $m$ push outcomes with small variations in distance and angle, weighted by a 2D Gaussian centered on that normal (so outcomes closer to that vector receive higher likelihood). Each hypothesized outcome is evaluated on $M$ by checking whether the displaced object would overlap with other objects, and classified into four categories: safe, low-cost contact, high-cost contact or contact with the target. Push outcomes in the same category are assigned an identical safety score $u\in(0,1]$, with larger values indicating safer outcomes. The aggregated safety score at $(x,y)$ is:
$f_{s}(x,y)=\frac{\sum_{i=1}^{m}l_{i}u_{i}}{\sum_{i=1}^{m}l_{i}},$ (1)
where $u_{i}$ encodes the safety level, and $l_{i}$ is the likelihood of the $i$-th push outcome. This score is then used to update the cost map at $(x,y)$, creating the final anisotropic cost map $M^{\prime}$ where each point has an associated directional push score:
$M^{\prime}[x,y]=\alpha M[x,y]+(1-\alpha)[10-10f_{s}(x,y)].$ (2)
Here, $\alpha\in(0,1)$ is a weight parameter to balance the object cost and the directional safety score $f_{s}\in[0,1]$, and the constant 10 ensures the updated cost in $M^{\prime}$ remains consistent with the original object cost range in $M$.
IV-B2 Contact-Aware A* Planning
To generate a plan, we search for a sequence of primitives that guides the end-effector to the target while accounting for directional costs in $M^{\prime}$. We formulate this as a graph search problem and use A* [11], which is well-suited for this setting and provides deterministic control compared to sampling-based algorithms. The planner’s state is defined as $S=(p,r,\mathcal{D})$, where $(p,r)$ is the robot end-effector position and orientation in 2D space and $\mathcal{D}$ is a list that tracks the cumulative 2D displacement of low-cost objects. At each step, the planner expands nodes using the three primitives. For Push, we estimate the object displacement using $d=D\cdot\big(\cos(\varphi)\big)^{\gamma}$, where $D$ is the maximum displacement when pushing in the optimal direction, $\varphi$ is the angular deviation between the push direction and the object’s reverse surface normal at the contact point, and the decay parameter $\gamma$ is set to $1.3$. IMPACT extends standard A* search by incorporating world state changes caused by pushes. Each valid Push creates a new branch in the search tree where an object has been permanently displaced to a new position. The resulting state, $S^{\prime}=(p^{\prime},r^{\prime},\mathcal{D^{\prime}})$, contains the updated robot pose and an updated world configuration.
A* search evaluates each state $s$ using the cost function $f(s)=g(s)+h(s)$, where $g(s)$ is the accumulated path cost from the start to state $s$, and $h(s)$ is a heuristic estimate of the remaining cost to the goal. We use the standard Euclidean distance to the target object as the heuristic $h(s)$. The accumulated path cost $g(s^{\prime})$ is updated at each step with action $a:s=(p,r,\mathcal{D})\rightarrow s^{\prime}=(p^{\prime},r^{\prime},\mathcal{D^{\prime}})$ as:
$g(s^{\prime})=g(s)+C(a)+P(s^{\prime}),$ (3)
where $C(a)$ is the action cost and $P(s^{\prime})$ is a gripper placement penalty. We define action costs based on the primitive. Move assigns a cost proportional to the Euclidean distance between two end-effector positions. Rotate applies a fixed cost for orientation changes. Push generates its cost from the value $M^{\prime}[x,y]$ at the contact point $(x,y)$. The penalty $P(s^{\prime})$ consists of three terms: (i) the average cost of the region the gripper overlaps, (ii) a collision penalty if the gripper collides with an object, and (iii) a proximity penalty proportional to the distance to high-cost objects. We defer additional details to the Appendix. By combining these costs, the planner can find a trajectory that minimizes risk while permitting acceptable contact when necessary; see Fig. 3 for more information about the A* planner. The path is then converted back to 3D space for evaluation and execution.
Figure 3: A* planner decision-making in two key scenarios, where the books and pack of chips are low-cost objects, while the mug and the stack of bowls are high-cost objects. The colored border around low cost objects visualizes anisotropic costs (red is unsafe, green is safe). On the left, the planner avoids a direct Push towards the stack of bowls (cost $\infty$) and instead chooses a low-cost Rotate to navigate between objects (cost $7.0$). On the right, by planning several steps ahead, it finds an efficient path by rotating and pushing stack of books (total cost $23.5$). It avoids a simpler but high-cost detour that only considers Move (cost $50.0$).